### v1

# A concise 40 T pulse magnet for condensed matter experiments

Abstract
There is a growing interest in using pulsed high magnetic field as a controlling parameter of physical phenomena in various scientific disciplines, such as condensed matter physics, particle physics, plasma physics, chemistry and biological studies. We devised a concise and portable pulsed magnetic field generator that produces a 40 T field with a pulse duration of 2 ms. It is assembled using only off-the-shelf components and a homemade coil that leverages small computers, Raspberry Pi, and Python codes. It allows for straightforward modification for general purposes. As working examples, we show representative applications in condensed matter experiments of magnetoresistance, magnetization, and magnetostriction measurements for graphite, NdNi2P2, and NdCo2P2, respectively, with the maximum magnetic field of 41 T and the lowest temperature of 4.2 K.
A magnetic field is a critical control parameter in condensed matters that directly affects electron spin, electron’s orbital angular momentum, the phase of electron’s wavefunction, and nuclear spins in materials [1, 2, 3, 4]. Not only that, researchers in diverse disciplines are investigating the magnetic field effects on vacuum fields [5], electromagnetic fluid dynamics in astrophysics [6, 7, 8], high-power laser plasma [9], electrocatalytic reaction [10], the magnetoreceptivity of songbirds [11, 12, 13], and magnetotactic bacteria [14, 15, 16]. The higher the magnetic fields get, the more evident the effect becomes. In these ways, the demand for pulsed magnetic fields is growing in various scientific fields.
Obtaining high magnetic fields is technically demanding [17]. Superconducting magnets in laboratories provide DC magnetic fields of 20 T. DC magnetic fields up to 45 T and pulsed high magnetic fields beyond 50 T are usually obtained only in large facilities for high magnetic fields. Besides, localizing the magnets for each experiment is challenging, which poses a barrier to implementing high magnetic field experiments. On the other hand, a pulsed magnet up to 40 T is obtained casually. Small-scale pulse magnets can generate a maximum field of 50 T with a 100 times smaller pulse power system than those in large facilities. Small pulse magnets have been proven quite compatible with quantum beams for the research of magnetic materials [18, 19, 20, 21] on vacuum and plasma [22, 23]. This technique has been mainly used for beam science because quantum beams require a relatively small magnetic field volume.
Figure 1: (a) Drawing of the main circuit. (b) Schematic image of the assembled pulse power system. (c) The schematic drawing of a pulse magnet.
Figure 2: (a) Measured magnetic field profile as functions of time generated using the mini-pulse power and the mini-coil. (b) The relation of charging voltage and the maximum current. (c) The relation of charging voltage and the maximum magnetic field. (d) The relation of maximum current and maximum magnetic field.
Figure 3: (a) A cross-section view of the He flow cryostat and the mini-coil. (b) Oblique view of the He flow cryostat and the mini-coil. (c) Magnification of the cross-section view of the cryostat and the mini-coil.
Small pulse magnets are even more promising for investing in broader scientific disciplines, including macroscopic laboratory experiments. However, such reports are missing. Here, we show a portable pulse magnet system that weighs 50 kg in total, generating 40 T powered by RaspberryPi and Python codes. It allows nearly everybody to have their own 40 T magnet in their laboratory, enabling rapid implementation of extremely high magnetic fields to new scientific fields. As a working example, we show its application to condensed matter experiments with a good data quality.
Figure 4: (a) Waveform of the pulsed magnetic field. (b) Magnetoresistance of HOPG as a function of time. The inset on the left shows the magnification of the same plot where the CDW transition occurs. The inset on the right shows a photo of the HOPG sample with the 4 contacts for the 4 probe measurement. (c) Magnetoresistance of HOPG as a function of the magnetic field. The inset on the left shows a schematic drawing of the 4 probe measurement and orientations of the magnetic field, current, and crystal axis. The inset on the right shows the magnification of the same plot where the CDW transition occurs.
Figure 5: (a) Waveform of the pulsed magnetic field. (b) Measured signal of induction by $dM/dt$ of NdNi2P2. The inset shows a photo of the inductive magnetization pick-up coil. (c) Magnetization of NdNi2P2 as a function of the magnetic field.
Figure 6: (a) Waveform of the pulsed magnetic field, (b) Magnetostriction of NdCo2P2 as a function of time. The inset shows a schematic drawing of magnetostriction measurement using FBG glued to the sample, (c) Magnetostriction of NdCo2P2 as a function of the magnetic field. The inset shows a photo of the FBG and sample glued with SK-229 (Nitto).
The primary circuit involves three parallel capacitors (800 $\mu$F, 2000 V) to form $C=2.4$ mF (EPCOS: TDK Lambda), five times lighter than the conventional oil film capacitor. The pulsed current is charged and discharged by a high-voltage charger (HUNS 3P117: Matsusada Precision), thyristor (5STP 17H5200: Hitachi ABB), Dump resistor, and high-voltage relay (Relay E12-NC: ROSS Engineering). The thyristor’s trigger circuit consists of a hand-made pulse transformer, optocoupler (FOD3182: Onsemi), and power MOSFET (2SK4017: Toshiba), which electrically separates external triggers from the high-voltage circuit. The high voltage charger is controlled with variable DC voltage output (I2C type DC10V output: NCD). The high-voltage relay is controlled with a small relay array (Bit-trade one). The charging voltage is monitored using an Analogue-Digital converter (I2C type ADS1015: Analog devices). All components are controlled by Python codes with Raspberry Pi, where the detailed Python codes are available in the Supplementary Note [24]. A mini-coil is made from a round wire with a diameter $\phi$1 mm Cu-Ag alloy (Ag 10 % in weight) and polyamide-imide enamel insulation. We used a semi-auto coil winder made of a speed-control and torque-control motor. The bore, winding number, and layers are 8 mm, 14, and 10, respectively. The drawing of the circuit, image of the assembled system, and coils are shown in Fig. 1.
With the mono-coil, we obtained the highest magnetic field of 40.7 T at a current of 5.6 kA with a charging voltage of 2100 V as shown in Fig. 2(a), where the initial temperature of the coil is 78 K immersed in liquid N2 bath. As shown in Figs. 2(b) and 2(c), $I_{\rm{max}}$ and $B_{\rm{max}}$ as a function of charging voltage shows a deviation from the linear relation above 1800 V, indicating that the coil heats up by Joule heating resulting in higher resistance. The result is in agreement with the result of the numerical calculation.
As a working example of the pulsed high magnetic fields, we show the magnetoresistance, magnetization, and magnetostriction measurements for highly oriented pure graphite (HOPG), NdNi2P2, and NdCo2P2, respectively, in our laboratory. To implement the low-temperature environment, we devised a plastic vacuum chamber and a plastic He flows cryostat [25] as shown in Fig. 3. First, we show the magnetoresistance measurement of HOPG at 4.5 K. The electric conductivity is measured using the four-probe method with an AC excitation at 1 MHz and 1 V. The voltage at the sample is amplified using a pre-amplifier (SR445: Stanford research) and recorded with a digital oscilloscope (PicoScope 4284A). The magnetoresistance of the sample is obtained by post-processing the recorded AC voltage of the sample using the numerical lock-in method [26]. The giant change in magnetoresistance with a saturating behavior is observed as shown in Figs. 4(b) and 4(c). The kink features at 35.1 T, which is magnified in the insets of Figs. 4(b) and 4(c), indicate a charge density wave (CDW) transition, being in good agreement with previous reports [27, 28].
Next, we show a magnetization measurement of a ferromagnetic metal NdNi2P2 at 4.2 K up to 35.1 T, the moment of which is dominated by the local moment of Nd [29]. For measurement, the powder sample is fused in a Kapton tube installed in a parallel-type induction coil [30]. We observe the ferromagnetic moment as the initial ramp of magnetization followed by a gradual increase of magnetization with increasing magnetic field, as shown in Fig. 5(b) and 5(c). The gradual increase is observed with high resolution, as shown in the magnified pink line in Fig. 5(b). The observed open loop is not an intrinsic hysteresis of the sample but an error in the measurement.
Lastly, we show the result of magnetostriction of an antiferromagnet NdCo2P2 [31, 32] at 4.2 K in Fig. 6(b) and 6(c). The magnetostriction measurement uses fiber Bragg grating and optical filter method [33, 34]. The results show a linear increase of magnetostriction up to 35.1 T, showing that a higher magnetic field is necessary for the magnetic saturation of the Nd moment. This indicates a relatively strong antiferromagnetic coupling between the Nd moment.
In summary, we show a pulse magnet with a simple design that weighs only 50 kg. It is easy to use and modify using Raspberry Pi and Python, whose detail is presented in the supplemental material [24]. The application to condensed matter measurements is shown with good-quality data. The concise pulse magnet system can provide broad scientific disciplines with pulsed high magnetic fields, such as particle physics, plasma physics, chemistry, and biology. We are also working on novel pulse magnets, such as a vector pulse magnet and a repetitive pulsed high magnetic field.
Acknowledgements.
Discussions with Y. H. Matsuda and H. Nojiri are acknowledged. This work is supported by MEXT LEADER program No. JPMXS0320210021, JST FOREST program No. JPMJFR [^2] W, JSPS Grants-in-Aid for Scientific Research No. 23H04861, 23H04859, 23H01121.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
DATA AVAILABILITY
The data supporting this study’s findings are available from the corresponding author upon reasonable request.

### v2

@@
 As working examples, we show representative applications in condensed matter experiments of magnetoresistance, magnetization, and magnetostriction measurements for graphite, NdNi2P2, and NdCo2P2, respectively, with the maximum magnetic field of 41 T and the lowest temperature of 4.2 K.
+I Introduction
 A magnetic field is a critical control parameter in condensed matters that directly affects electron spin, electron’s orbital angular momentum, the phase of electron’s wavefunction, and nuclear spins in materials [1, 2, 3, 4].
@@
 Small-scale pulse magnets can generate a maximum field of 50 T with a 100 times smaller pulse power system than those in large facilities.
-Small pulse magnets have been proven quite compatible with quantum beams for the research of magnetic materials [18, 19, 20, 21] on vacuum and plasma [22, 23].
+Small pulse magnets have been proven useful for the research of magnetic materials using x-ray diffractions [18, 19, 20, 21], x-ray spectroscopy [22, 23, 24, 25], neutron scatterings [26, 27, 28].
+It has also been applied for the investigation of magnetic field effects on vacuum [29] and plasma [30].
 This technique has been mainly used for beam science because quantum beams require a relatively small magnetic field volume.
-Figure 1: (a) Drawing of the main circuit. (b) Schematic image of the assembled pulse power system. (c) The schematic drawing of a pulse magnet.
+Figure 1: (a) Sketch of the pulse power circuit.
+(b) Schematics of the assembled pulse power system.
+(c) The schematic drawing of a pulse magnet.
+(d) A photo of our compact pulse-magnet-lab.
+with 3 assembled pulse powers.
+The inset is a photo of a homemade coil.
 Figure 2: (a) Measured magnetic field profile as functions of time generated using the mini-pulse power and the mini-coil.
@@
 (c) Magnification of the cross-section view of the cryostat and the mini-coil.
+(d) A photograph of the 3D printed vacuum chamber and the FRP cryostat.
 Small pulse magnets are even more promising for investing in broader scientific disciplines, including macroscopic laboratory experiments.
@@
 The inset on the right shows the magnification of the same plot where the CDW transition occurs.
-Figure 5: (a) Waveform of the pulsed magnetic field.
-(b) Measured signal of induction by $dM/dt$ of NdNi2P2.
+(d) Waveform of the pulsed magnetic field.
+(e) Measured signal of induction by $dM/dt$ of NdNi2P2.
 The inset shows a photo of the inductive magnetization pick-up coil.
-(c) Magnetization of NdNi2P2 as a function of the magnetic field.
-Figure 6: (a) Waveform of the pulsed magnetic field, (b) Magnetostriction of NdCo2P2 as a function of time.
-The inset shows a schematic drawing of magnetostriction measurement using FBG glued to the sample, (c) Magnetostriction of NdCo2P2 as a function of the magnetic field.
+(f) Magnetization of NdNi2P2 as a function of the magnetic field.
+(g) Waveform of the pulsed magnetic field, (h) Magnetostriction of NdCo2P2 as a function of time.
+The inset shows a schematic drawing of magnetostriction measurement using FBG glued to the sample, (i) Magnetostriction of NdCo2P2 as a function of the magnetic field.
 The inset shows a photo of the FBG and sample glued with SK-229 (Nitto).
+II Method
 The primary circuit involves three parallel capacitors (800 $\mu$F, 2000 V) to form $C=2.4$ mF (EPCOS: TDK Lambda), five times lighter than the conventional oil film capacitor.
@@
 The result is in agreement with the result of the numerical calculation.
+III Results
 As a working example of the pulsed high magnetic fields, we show the magnetoresistance, magnetization, and magnetostriction measurements for highly oriented pure graphite (HOPG), NdNi2P2, and NdCo2P2, respectively, in our laboratory.
@@
 3.
+Using this He flow cryostat, the sample temperature is variable from 300 K to 4.2 K by regulating the He gas flow.
 First, we show the magnetoresistance measurement of HOPG at 4.5 K.
-The electric conductivity is measured using the four-probe method with an AC excitation at 1 MHz and 1 V.
+The electric conductivity is measured using the four-probe method with an AC excitation at 1 MHz and 1 mA.
 The voltage at the sample is amplified using a pre-amplifier (SR445: Stanford research) and recorded with a digital oscilloscope (PicoScope 4284A).
@@
 We observe the ferromagnetic moment as the initial ramp of magnetization followed by a gradual increase of magnetization with increasing magnetic field, as shown in Fig.
-5(b) and 5(c).
+4(e) and 4(f).
 The gradual increase is observed with high resolution, as shown in the magnified pink line in Fig.
-5(b).
+4(e).
 The observed open loop is not an intrinsic hysteresis of the sample but an error in the measurement.
 Lastly, we show the result of magnetostriction of an antiferromagnet NdCo2P2 [38, 39] at 4.2 K in Fig.
-6(b) and 6(c).
+4(h) and 4(i).
 The magnetostriction measurement uses fiber Bragg grating and optical filter method [40, 41].
@@
 This indicates a relatively strong antiferromagnetic coupling between the Nd moment.
+IV Summary
 In summary, we show a pulse magnet with a simple design that weighs only 50 kg.
@@
 The concise pulse magnet system can provide broad scientific disciplines with pulsed high magnetic fields, such as particle physics, plasma physics, chemistry, and biology.
-We are also working on novel pulse magnets, such as a vector pulse magnet and a repetitive pulsed high magnetic field.
+We are also working on novel pulse magnets, such as a vector pulse magnet and AC 100 Hz pulsed high magnetic field.
+Supplemental Material
+Supplementary Material shows how to build a 40 T generator, including the capacitor bank and a mini-coil.
 Acknowledgements.
@@
 JPMXS0320210021, JST FOREST program No.
-JPMJFR [^2] W, JSPS Grants-in-Aid for Scientific Research No.
-23H04861, 23H04859, 23H01121.
+JPMJFR [^2] W, JSPS Grant- in-Aid for Scientific Research on Innovative Areas (A) 23H04861, 23H04859, 43 and Grant-in-Aid for Scientific Research (B) 23H01121.
 AUTHOR DECLARATIONS